Forty eight grams of carbon is added and reacted to 134grams of oxygen to produce Carbon dioxide. Find (a) the balance chemical equation (4points), (b) the mass of the product (4points), (c) the limiting reactant (4points), (d) the mass of excess reactant not used in the reaction (6points). Write the complete formula and solution for each problem.You can use a separate sheet if needed. Use whole numbers only.

1
Expert's answer
2022-03-25T11:10:05-0400

(a) C + O2 = CO2

(c) n(C) = m(C)/Mr(C) = 48 g / 12 g/mol = 4 mol

n(O2) = m(O2)/Mr(O2) = 134 g / 32 g/mol = 4.1875 mol

n(C) < n(O2)

C is limiting reactant

(b) n(C) = n(CO2) = 4 mol

m(CO2) = n(CO2) × Mr(CO2) = 4 mol × 44 g/mol = 176 g

(d) O2 is excess reactant

n(O2)not reacted = 4.1875 mol - 4.0 mol = 0.1875 mol

m(O2)not reacted = n(O2)not reacted × Mr(O2) = 0.1875 mol × 32 g/mol = 6 g